OpenStudy (anonymous):

Find the remainder when f(x) = 2x3 – 12x2 + 11x + 2 is divided by x – 5. Answer: A) 3 B) –7 C) 7 D) –3 Could someone help me please:)

OpenStudy (anonymous):

find \[f(5)\]and that is your answer

OpenStudy (ash2326):

let's try to create x-5 in each of the terms, then we'll take it common and the remaining part will be the remainder f(x)=2x^3-12x^2+11x+2 writng 2x^3 as----> 2x^2(x-5) this had added an extra term -10x^2, so we'll add a 10x^2 f(x)=2x^2(x-5)+10x^2-12x^2+11x+2 f(x)=2x^2(x-5)-2x^2+11x+2 =2x^2(x-5)-2x(x-5)-10x+11x+2 ( we wote -2x^2 as -2x(x-5) this added a 10 x , so we' have added -10x) =2x^2(x-5)-2x(x-5)+x+2 =2x^2(x-5)-2x(x-5)+x-5+3 =(x-5)(2x^2-2x+1) +3 =so the quotient is 2x^2-2x+1 and the remainder is 3 option A

OpenStudy (anonymous):

Thank you ashley!!!!

OpenStudy (anonymous):

*Ash:)

OpenStudy (anonymous):

Actually its not option A; got it wrong

OpenStudy (anonymous):

Since the remainder is 7, the answer would be C

OpenStudy (anonymous):

Im in flvs too btw :3